# Specific heat of Gd3+ and Eu2+-based magnetic compounds

### v3

@@ IV Effect of anisotropy
IV.1 Mean-field Approximation
@@
IV.2 Beyond Mean-field
Figure 4: Specific heat $C/T$ versus $T/T_{C}$ for the EuPdSn2(16) structure with varying axial anisotropy $K$ ($J_{\delta}=-1$). The critical temperatures for K=0.6,0.3,0,0.3,0.6K=-0.6,-0.3,0,0.3,0.6 are approximately $78,74,68.5,71.5,73$ respectively. To compare with Figure 3 the ratio $K/{k_{B}T_{c}}$ for these cases are .0077-.0077, 0.0040-0.0040,0,$0.0042$ and $0.0082$ respectively.
In this subsection, we discuss the effect of anisotropy in cases where the fluctuations of the effective magnetic field are also included. Figure 4 shows the specific heat for the EuPdSn2(16) structure (described in section VI) with varying anisotropy $K$ for moderate values ($|K|\lesssim J_{\delta}$). For $K=0$, the system is isotropic, while negative $K$ values introduce axial anisotropy, favoring alignment along the $z$-axis. $K>0$ introduces a plane anisotropy favoring alignment in the $xy$-plane. Qualitatively, the results for $T<T_{C}$ mirror those of the mean-field approximation, with anisotropy shifting entropy from the region $T<0.5T_{C}$ to $0.5T_{C}<T<T_{C}$, and the effect is more pronounced for negative $K$.
@@ IV.2 Beyond Mean-field
V Comparison of Different Structures with the Same Effective Number of Neighbors
@@
Figure 5: Specific heat $C/T$ versus $T/T_{C}$ for various lattice structures with different values of effective neighbor count $z$ (defined by Eq. 4), with $J_{1}=-1$. The critical temperatures $T_{C}$ and $z$ values are tabulated in Table 3. Top panel: Diamond lattice with varying $J_{2}$ interactions ($z=4,6,8,10,12,14,16$ from bottom to top). Inset: $z$ as a function of $J_{2}$. Middle panel: Structures with $z=8$ or $12$ (diamond, BCC and FCC). Bottom panel: Structures with $z=6$ or $18$ (diamond, SC, BCC and FCC).
@@
Our central finding is that the specific heat of magnetic systems is primarily determined by the effective number of neighbors $z$, rather than the specific lattice structure or interaction pattern. In Section III, we analyzed the impact of varying the number of neighbors $z$ on the specific heat, while keeping $J_{\delta}=J_{1}$. Here, we extend this analysis by investigating how changes in $J_{\delta}$ affect the specific heat, while maintaining the effective number of neighbors $z$ constant [see Eq. (4)]. The parameter $z$ quantifies the relative strength of spatial and quantum fluctuations. Physically, $z$ represents the effective number of neighbors that meaningfully contribute to the exchange field experienced by each spin, accounting for both geometric coordination and relative interaction strengths.
@@
To illustrate this principle, we use a diamond lattice with interactions $J_{1}=-1$ and varying $J_{2}$ as a guide. This allows us to explore cases where $z$ varies between $4$ (just nearest neighbors) and $16$ (nearest and next-nearest neighbors). In Figure 5(a), we show how the specific heat changes with $J_{2}$, with $z$ values ranging from $4$ to $16$. The specific heat curves closely follow the results shown in Figure 2, with even the $z=8$ and $z=12$ cases being nearly degenerate.
In the middle panel (b) and bottom panel (c) of Figure 5, we compare these diamond and non-uniform BCC and FCC results against lattices with uniform couplings (SC(6), BCC(8), FCC(12), FCC(18)). The specific heat curves for diamond and uniform coupling lattices with the same $z$ are strikingly similar. Also the diamond lattice with $J_{2}=-0.155$ ($z=8$) matches the BCC lattice with $J_{2}=-8.0$ ($z=8$), and the diamond lattice with $J_{2}=-0.333$ ($z=12$) matches the FCC lattice with $J_{2}=-4.0$ ($z=12$) . Despite the drastic changes in interaction magnitudes, the specific heat curves align closely, underscoring the dominance of $z$ over detailed interaction strengths. Even for the FCC lattice with $J_{2}=-4$, where curves do not overlap, the comparison remains good.
V.1 Getting information beyond mean field
@@
The key insight of our work is that the effective number of neighbors $z$ (defined in Eq. 4) is the dominant parameter determining the specific heat, rather than the specific lattice structure or interaction pattern. This allows us to extract valuable information about the underlying exchange interactions that goes beyond what can be obtained from mean-field theory alone.
@@
To illustrate this, consider the diamond lattice with varying $J_{2}$ (Fig. 5). For different values of $J_{2}$, we obtain different effective number of neighbors $z$. Crucially, when we compare diamond lattices with different $J_{2}$ values but the same $z$ to reference lattices with uniform couplings (SC, BCC, FCC), we find that the specific heat curves match almost perfectly. This demonstrates that the specific heat is insensitive to the detailed interaction pattern as long as $z$ is the same.
@@
Let’s examine how this works in practice using the examples from Fig. 5:
@@
    For the diamond lattice with $J_{2}=-0.080$ ($z=6.03$), the specific heat matches SC(6). Using Eq. 4 with $z_{1}=4$ and $z_{2}=12$ for the diamond lattice, we can solve for $J_{2}/J_{1}$:
    $1/z=\frac{z_{1}J_{1}^{2}+z_{2}J_{2}^{2}}{(z_{1}J_{1}+z_{2}J_{2})^{2}}=1/6.$
    With $J_{1}=-1$, this yields $J_{2}/J_{1}\approx 0.0787$, which is very close to the actual value of $0.080$. This demonstrates that the effective $z$ parameter can be used to extract the ratio of interaction strengths with high accuracy. Using the critical temperature as an energy scale, we can even determine the absolute values of $J_{1}$ and $J_{2}$.
@@
    Diamond lattices with $J_{2}=-0.155$ ($z=8$) and $J_{2}=-0.333$ ($z=12$) both match BCC(8) and FCC(12) as specific heat with $z=8$ and $12$ are very close. This indicates that these lattices require next-nearest-neighbor interactions for an accurate description. The analysis limits the possible parameter sets to $J_{2}/J_{1}\approx 0.155$ or $J_{2}/J_{1}\approx 0.333$, providing valuable constraints on the interaction strengths.
@@
    As in the previous case BCC($J_{2}=-0.40$) and BCC($J_{2}=-8.0$) lead to multiple possibilities: if $z=8$, then $J_{2}/J_{1}=0$ or $8$; if $z=12$, then $J_{2}/J_{1}=0.390$ or $2.276$.
@@
    For FCC($J_{2}=-4.0$), only two possibilities exist: $J_{2}/J_{1}=4$ ($z=12$) or $J_{2}/J_{1}=12.3$ ($z=8$).
  5. 5.
    For BCC($J_{1}=J_{2}=-1.0$, $J_{3}=-0.175$), the specific heat matches FCC(18) and lattices with $z=16$. This indicates that both nearest-neighbor (8) and next-nearest-neighbor (6) interactions are relevant. Considering interactions up to third order leads to a continuum of possibilities that can be further constrained by additional data.
@@
The practical significance of this finding is substantial. By comparing experimental specific heat to our reference curves, one can extract the effective number of neighbors $z$ and, from $z$, infer the relative strengths of different exchange interactions. This information is inaccessible through mean-field analysis alone and can guide the development of more accurate theoretical models. For example, if a specific heat of a SC compound matches our reference curve for $z=8$ (and so it also matches $z=12$), we can immediately infer that the system likely has a combination of nearest-neighbor and next-nearest-neighbor interactions with a ratio $J_{2}/J_{1}\approx 0.08$ ($z=8$) or $J_{2}/J_{1}\approx 0.25$ ($z=12$).
This framework represents a significant advance in our ability to interpret magnetic specific heat data for 4f7 systems, as demonstrated in Section VI where we apply these principles to fit experimental data for six different compounds.
VI Fits of experimental data
In this section, we present fits of the specific heat of six compounds: GdNi3Ga9, GdCu2Ge2, GdPbBi, GdNiSi3, Eu2Pd3Sn3, and EuPdSn2 (see Figure 6). The fitting procedure uses the experimentally determined critical temperature $T_{N}$ as a fixed input, with the effective number of neighbors $z$ and axial anisotropy $K$ as free parameters. For Gd compounds, $K=0$ was used as the anisotropy term was found to be negligible. The assumption of uniform exchange interactions ($J_{\delta}=J$) is justified by our theoretical results showing that different $J_{\delta}$ values with the same effective number of neighbors $z$ yield nearly identical specific heat curves (see Section V).
@@ VI Fits of experimental data
Table 4: Parameters used in the fits: $T_{N}$ is the Neel temperature, $z_{0}$ is the number of nearest neighbors in the reported structure (compatible with 3D connectivity), $z$ (fit) is the effective number of neighbors, $J$ is the interaction strength, and $K$ is the anisotropy. This value correspond to $J_{1}=J_{2}$ and $J_{3}=0.27J_{1}$, see main text.
GdNi3Ga9 [33, 34] exhibits a layered honeycomb structure with $z_{0}=4$ (three intralayer and one interlayer nearest neighbors), consistent with $z=4$ found in the fit. The fit (Figure 6a) reproduces the experimental specific heat curve with excellent accuracy, with the peak at $T_{C}=20$ K matching the experimental value. The small $z$ value indicates significant spatial and quantum fluctuations, which is consistent with the observed broadening of the specific heat peak. As LHc(4,AF) lattice has $T_{N}/{|J_{1}|}\sim 13$ while experimentally $T_{N}=20K$ we can deduce $J\sim 20$$13\sim 1.5K$ corresponding to an antiferromagnetic interaction.
For the rest of the structures, ferromagnetic interactions are assumed. As stated in Section III, as long as there is no significant frustration, the difference between ferromagnetic and antiferromagnetic interactions for $z>6$ is only noticeable at very low temperatures, well below the temperature range of interest in this study. In figure 5 we show the specific heat of a BCC lattice with AF interactions that compares well with the FM case. The key features of the specific heat curves, such as the shape of the peak and the high-temperature tail, are primarily determined by the effective number of neighbors and the single-ion anisotropy. These features are largely insensitive to the sign of the exchange interaction in the absence of strong frustration. For this reason we compare the experimental Neel temperature ($T_{N}$) with the theoretical Curie temperature ($T_{C}$).
@@
  • $J_{3}/J_{1}=3.7$ ($z=12$) which we can disregard has it seems unphysical to have a larger strength associated to a more distant bond.
  • $J_{3}/J_{1}=0.27$ ($z=12$) appears to be the only sensitive solution.
Without the actual simulation for this combination of couplings, we can only estimate the coupling’s strength. To get the strength of the coupling we use the MF result for the critical temperature (see section IV.1)
@@
from which we obtain $|J_{1}|=|J_{2}|\sim 0.20$K and $|J_{3}|\sim 0.05$K. From this results we can not determine the signs of the coupling. But this antiferromagnetic compound at least one of these coupling should be antiferromagnetic.
GdPdBi [36] is a half-Heusler structure. Gd sublattice in this compound forms a FCC structure. We obtain a very good fit with the LHc(10) lattice. Using $z=10$ leads to $J2/J_{1}\sim 6.1$ which seems unphysical. As FCC(12) ($T_{C}/{|J_{1}|}\sim 50$) also accounts for the experimental data we use that lattice to estimate the interaction strength. As in the previous compound we suppose that a NN interaction alone scenario is closer to the compound. Experimentally $T_{N}=13$K so we deduce $|J|\sim 13$K/50=0.26/50=0.26K. For this antiferromagnetic compound an antiferromagnetic coupling is expected.
The crystal structure of GdNiSi3 is an orthorhombic system. Gd atoms forms a bilayered arrangement of Gd. Only interaction at the 7th nearest neighbor connects bilayers to provide a tridimensional connectivity ($z_{0}=17$). A good description of the experimental data is obtain considering up to 10th nearest-neighbor ($z=25$) while the best fit is seen considering up to 13th NN ($z=33$). This suggests that long range interactions, well beyond first neighbors, are relevant in this compound. Given the complexity of these interactions we only consider the uniform $J_{\delta}$ case. The GdNiSi3(33) lattice has $T_{C}/{|J_{1}|}\sim 145$ while experimentally $T_{N}=22$K so we can deduce $|J|\sim 22$K/145=0.15/145=0.15K. The large amount of couplings needed (up to 13 nearest-neighbors) certainly come from a distribution of interactions. This distribution could be determined with first principle calculations that considers at least this coordination.
@@
Now turning to Eu compounds, the importance of $K$ for the Eu compounds may be attributed to the $\sim 25\%$ larger ionic radius of Eu2+ compared to Gd3+ [37], which makes the former more susceptible to crystal-field effects.
@@
The Eu sublattice of Eu2Pd3Sn can be though as a stacking of corrugated hexagonal layers in the ab plane. The smaller coordination number to have a tridimensional connectivity is $z_{0}=4$. This involves up to third order nearest-neighbor interactions. To properly compare the specific heat of Eu2Pd3Sn [38] with a Hamiltonian model we have to consider up to 6th nearest-neighbor in this lattice leading to a coordination $z=14$. The point group at the Eu sites is $C_{1h}$, which allows for the inclusion of the anisotropy term $H_{A}$, as discussed in section I.2. So an anisotropy term with $K/|J_{1}|=-0.4$ was included. With this anisotropy the Eu2Pd3Sn(14) lattice has $T_{C}/{|J_{0}|}\sim 66$. Experimentally $T_{N}=47$K so we can deduce, assuming uniform couplings, $|J|\sim 47$K/66=0.71/66=0.71K.
@@
Finally, the Eu sublattice of EuPdSn2 can be though as a stacking of rectangular layers in the ac plane. The smaller coordination number to have a tridimensional connectivity is $z_{0}=8$. This involves up to 3 nearest-neighbor interactions. To properly compare the specific heat of EuPdSn2 [29] with a Hamiltonian model we have to consider up to 7th nearest-neighbor in this lattice leading to a coordination $z=22$. The point group at the Eu sites is $C_{2v}$, which also allows for the inclusion of the term $H_{A}$. So an anisotropy term with $K/|J_{0}|=-0.9$ was included. With this anisotropy the EuPdSn2 lattice has $T_{C}/{|J_{0}|}\sim 112$. Experimentally $T_{N}=12$K from which we deduce $|J|\sim 12$K/112=0.11/112=0.11K. The discrepancies near $T_{N}$ likely originate from factors not explicitly included in our model, such as more complex magnetic ordering (e.g., a collinear amplitude-modulated structure) resulting from modulated interactions and crystal-field effects, as discussed in Ref. 14.
@@
Since all the compounds contain transition-metal ions, which may be magnetic, one might question their influence on the specific heat. However, our analysis shows that the entropy obtained by integrating $C/T$ is close to $\ln(8)$ per rare-earth ion (consistent with $S=7/2$), indicating that any effects from transition metals are negligible.
@@
The success of these fits suggests that the effective number of neighbors $z$ is the dominant parameter determining the specific heat of 4f7 systems, with axial anisotropy $K$ providing additional refinement for Eu compounds. This approach significantly outperforms mean-field theory, which would fail to reproduce the observed specific heat curves (see Figure 2).
@@
Figure 6: Experimental specific heat $C/T$ versus $T/T_{N}$ for six 4f7 compounds (dots) and our theoretical fits (lines). $K=0$ for all Gd-based compounds (GdNi3Ga9, GdCu2Ge2, GdPbBi, GdNiSi3). Experimental data sources: Refs. 33, 34, 36, 35, 27, 29, 38.
VII Summary

### v4

@@ V Effect of anisotropy
V.1 Mean-field Approximation
@@
V.2 Beyond Mean-field
Figure 4: Specific heat $C/T$ versus $T/T_{C}$ for the EuPdSn2(16) structure with varying axial anisotropy $K$ ($J_{\delta}=-1$). The critical temperatures for K=0.6,0.3,0,0.3,0.6K=-0.6,-0.3,0,0.3,0.6 are approximately $78,74,68.5,71.5,73$ respectively. To compare with Fig. 3 the ratio $K/{k_{B}T_{c}}$ for these cases are .0077-.0077, 0.0040-0.0040,0,$0.0042$ and $0.0082$ respectively.
In this subsection, we discuss the effect of anisotropy in cases where the fluctuations of the effective magnetic field are also included. Fig. 4 shows the specific heat for the EuPdSn2(16) structure (described in section VII) with varying anisotropy $K$ for moderate values ($|K|\lesssim J_{\delta}$). For $K=0$, the system is isotropic. Negative $K$ values introduce axial anisotropy, favoring alignment along the $z$-axis, whereas positive $K$ values produce planar anisotropy, favoring alignment within the $xy$-plane. Qualitatively, the results for $T<T_{C}$ mirror those of the mean-field approximation, with anisotropy shifting entropy from the region $T<0.5T_{C}$ to $0.5T_{C}<T<T_{C}$, and the effect is more pronounced for negative $K$.
@@ V.2 Beyond Mean-field
VI Comparison of Different Structures with the Same Effective Number of Neighbors
@@
Figure 5: Specific heat $C/T$ versus $T/T_{C}$ for diamond and FCC lattices with different values of effective number of neighbors $z$ [defined by Eq. (4)], with $J_{1}=-1$. The critical temperatures $T_{C}$ and $z$ values are tabulated in Table 3. Top panel: Diamond lattice with varying interactions with $|J_{2}|\leq|J_{1}|$ ($z=4,6,8,10,12,14,16$ from bottom to top). Top panel inset: $z$ as a function of $J_{2}$ for diamond. Bottom panel: same results for FCC lattices with varying interactions with $|J_{2}|\geq|J_{1}|$ ($z=8,10,12,14,16,18$). Bottom panel inset: $z$ as a function of $J_{2}$ for FCC.
@@
Our central finding is that the specific heat of magnetic systems is primarily determined by the effective number of neighbors $z$, defined by Eq. (4) rather than the specific lattice structure or interaction pattern. The parameter $z$ quantifies the relative strength of spatial and quantum fluctuations. Physically, $z$ represents the effective number of neighbors that meaningfully contribute to the exchange field experienced by each spin, accounting for both geometric coordination and relative interaction strengths. In Section IV, we analyzed the impact of varying $z$ on the specific heat, while keeping $J_{\delta}=J_{1}$. Here, we extend this analysis by investigating how changes in $J_{\delta}$ affect the specific heat, while maintaining the effective number of neighbors $z$ constant [see Eq. (4)].
@@
To illustrate this principle, we use diamond and FCC lattices. These two systems allow us to explore cases where $z$ varies between 4 (just nearest neighbors) and 16 (nearest and next-nearest neighbors) for the diamond lattice, and between 8 and 18 for the FCC lattice. In Fig. 5(a) and (b), we show how the specific heat changes with $J_{2}$, for both lattices and for $z$ values ranging from 4 to 18. The insets in each panel display $z$ as a function of $J_{2}$ for the respective lattice. For the diamond lattice with $0\geq J_{2}\geq J_{1}$ ($|J_{2}|<|J_{1}|$), the specific heat curves closely follow the results shown in Fig. 2. For the FCC lattice with $|J_{2}|>|J_{1}|$, the specific heat appears almost independent of $J_{2}$. This behavior may be attributed to the FCC lattice’s non-bipartite nature, which could involve higher-order effective interactions that equilibrate fluctuations from first and second neighbors.
Figure 6: Specific heat $C/T$ versus $T/T_{C}$ for various lattice structures with different values of the effective number of neighbors $z$ [defined by Eq. (4)], with $J_{1}=-1$ and $0\geq J_{2}\geq J_{1}$. The critical temperatures $T_{C}$ and $z$ values are tabulated in Table 3. Purple lines correspond to diamond lattice results while black, brown, blue and green correspond to simple cubic, BCC, FCC and LHc respectively. Panels a) to f) correspond to $z$ from 8 to 16. Orange lines on panels a) and b) correspond to antiferromagnetic interactions.
A detailed comparison between diamond, simple cubic (SC), body-centered cubic (BCC), FCC, and layered honeycomb (LHc) lattices is shown in Figure 6. In this figure, we present results for $z$ values between 6 and 16 with $|J_{2}|\leq|J_{1}|$. The agreement of the specific heat for different lattices with varying $J_{2}$ but the same $z$ is excellent. We also show with orange lines the specific heat for non-frustrated antiferromagnetic cases with $z=6$ and $z=8$, showing a clear convergence toward the ferromagnetic results as $z$ increases.
Figure 7: Specific heat $C/T$ versus $T/T_{C}$ for various lattice structures with different values of effective number of neighbors $z$ [defined by Eq. (4)], with $J_{1}=-1$. The critical temperatures $T_{C}$ and $z$ values are tabulated in Table 3. Purple lines correspond to diamond lattice results while black, brown, blue and green correspond to simple cubic, BCC, FCC and LHc respectively. Panels a) to f) correspond to $z$ from 8 to 16. Orange lines on panels a) and b) correspond to antiferromagnetic interactions.
In Figure 7, we show the specific heat evolution when $|J_{2}|\geq|J_{1}|$. As a reference, the purple lines depict the specific heat for the diamond lattice with the corresponding $z$ value for $|J_{2}|\leq|J_{1}|$. We compare the diamond (green), BCC (brown), simple cubic (SC, black), and FCC (blue) lattices. For values of $|J_{2}|$ much larger than $|J_{1}|$ (panels a–c), there is a significant departure from the diamond reference. The fact that the FCC lattice (non-bipartite) shows the largest deviation compared to BCC and SC suggests that higher-order interactions could equilibrate fluctuations from first and second neighbors. In panel (f), we also show the BCC case with $z=18$ and three couplings, where the agreement with the SC and FCC lattices is remarkable.
These results demonstrate that the effective number of neighbors $z$ encapsulates the essential physics of spatial and quantum fluctuations for $0\geq J_{2}\geq J_{1}$ ($|J_{2}|\leq|J_{1}|$), and it provides a good approximation for $|J_{2}|>|J_{1}|$. This enables a universal description of specific heat across diverse magnetic structures, independent of their lattice geometry or interaction patterns.
VI.1 Getting information beyond mean field
@@
The key insight of our work is that the effective number of neighbors $z$ [defined in Eq. (4)] is the dominant parameter determining the specific heat, rather than the specific lattice structure or interaction pattern, when $0\geq J_{2}\geq J_{1}$. This allows us to extract valuable information about the underlying exchange interactions that goes beyond what can be obtained from mean-field theory alone.
@@
To illustrate this, consider the diamond lattice with varying $J_{2}$ (Fig. 6). For different values of $J_{2}$, we obtain different effective numbers of neighbors $z$. Crucially, when we compare diamond lattices with different $J_{2}$ values but the same $z$ to reference lattices with uniform couplings (SC, BCC, FCC, LHc), we find that the specific heat curves match almost perfectly. This demonstrates that the specific heat is insensitive to the detailed interaction pattern as long as $z$ is the same.
@@
Let’s examine how this works in practice using the examples from Fig. 6. Assume that an experimental specific heat curve coincides with one of those displayed in the figure, while the exchange interactions $J_{\delta}$ are initially unknown.
@@
    Diamond lattice with $J_{2}=-0.080$ ($z=6.03$): The specific heat matches SC(6). Using Eq. (4) with $z_{1}=4$ and $z_{2}=12$ for the diamond lattice, we solve for $|J_{2}/J_{1}|$:
    $1/z=\frac{z_{1}J_{1}^{2}+z_{2}J_{2}^{2}}{(z_{1}|J_{1}|+z_{2}|J_{2}|)^{2}}=1/6.$
    This yields $|J_{2}/J_{1}|\approx 0.0787$, very close to the actual value of $0.080$. This demonstrates that the parameter $z$ can be used to extract the ratio of interaction strengths with high accuracy. Using the critical temperature as an energy scale, we can even determine the absolute values of $J_{1}$ and $J_{2}$.
@@
    Diamond lattices with $J_{2}=-0.155$ ($z=8$) and $J_{2}=-0.333$ ($z=12$): Both match BCC(8) and FCC(12) as specific heat for these reference lattices are very close. This indicates that these diamond lattices require next-nearest-neighbor interactions for an accurate description. The analysis limits the possible parameter sets to $|J_{2}/J_{1}|\approx 0.155$ or $|J_{2}/J_{1}|\approx 0.333$, providing valuable constraints on the interaction strengths.
@@
    BCC lattice with $J_{2}=-0.40$: As in the previous case BCC($J_{2}=-0.40$) matches BCC(8) and FCC(12) and lead to multiple possibilities: if $z=8$, then $|J_{2}/J_{1}|=0$ or $8$; if $z=12$, then $|J_{2}/J_{1}|=0.390$ or $2.276$. This illustrates how $z$ can narrow down plausible interaction ratios, even when multiple solutions exist.
@@
    BCC lattice with $J_{1}=J_{2}=-1.0$, $J_{3}=-0.175$: The specific heat matches FCC(18). This indicates that at least both nearest-neighbor (8) and next-nearest-neighbor (6) interactions are relevant. Considering interactions up to third order leads to a continuum of possibilities that can be further constrained by additional data.
@@
The practical significance of this finding is substantial. By comparing experimental specific heat to our reference curves, one can extract the effective number of neighbors $z$ and, from $z$, infer the relative strengths of different exchange interactions. This information is inaccessible through mean-field analysis alone and can guide the development of more accurate theoretical models. For example, if the specific heat of a SC compound matches our reference curve for $z=8$ (thus it also matches $z=12$), we can immediately infer that the system likely has a combination of nearest-neighbor and next-nearest-neighbor interactions with a ratio $|J_{2}/J_{1}|\approx 0.08$ ($z=8$) or $|J_{2}/J_{1}|\approx 0.25$ ($z=12$).
This framework represents a significant advance in our ability to interpret magnetic specific heat data for 4f7 systems, as demonstrated in Section VII where we apply these principles to fit experimental data for six different compounds.
VII Fits of experimental data
In this section, we present fits of the specific heat of six compounds: GdNi3Ga9, GdCu2Ge2, GdPbBi, GdNiSi3, Eu2Pd3Sn3, and EuPdSn2 (see Figure 8). The fitting procedure uses the experimentally determined critical temperature $T_{N}$ as a fixed input, with the effective number of neighbors $z$ and axial anisotropy $K$ as free parameters. For Gd compounds, $K=0$ was used as the anisotropy term was found to be negligible. The assumption of uniform exchange interactions ($J_{\delta}=J$) is justified by our theoretical results showing that different $J_{\delta}$ values with the same effective number of neighbors $z$ yield nearly identical specific heat curves (see Section VI).
@@ VII Fits of experimental data
Table 4: Parameters used in the fits: $T_{N}$ is the Néel temperature, $z_{0}$ is the number of nearest neighbors in the reported structure (compatible with 3D connectivity), $z$ (fit) is the effective number of neighbors, $J$ is the interaction strength, and $K$ is the anisotropy. From the fit we can only conclude that $|J_{1}|<|J_{2}|$, see main text. This value correspond to $|J_{1}|=|J_{2}|$ and $|J_{3}|=0.05K$, see main text.
GdNi3Ga9 [34, 35] exhibits a layered honeycomb structure with $z_{0}=4$ (three intralayer and one interlayer nearest neighbors), consistent with $z=4$ found in the fit. The specific heat fit (Figure 8a) reproduces the experimental data with excellent accuracy. The low $z$ value indicates significant spatial and quantum fluctuations accounting for the observed broadened specific heat peak. As the LHc(4,AF) lattice (corresponding to an antiferromagnetic interaction) has $T_{N}/J_{1}\sim 13$ while experimentally $T_{N}=20K$ we can deduce $J\sim 20$$13\sim 1.5K$. This result of a dominant first-neighbor antiferromagnetic interaction is further supported by the observation of a small (1.8°) deviation from perfect AF alignment between magnetic moments in adjacent layers [35], which may arise from weak Dzyaloshinskii-Moriya interactions (DMI).
For the rest of the structures, ferromagnetic interactions are used for the fits. As stated in Section IV, as long as there is no significant frustration, the difference between ferromagnetic and antiferromagnetic interactions for $z>6$ is only noticeable at very low temperatures, well below the temperature range of interest in this study. In Fig. 6 we show the specific heat of a SC and a BCC lattice with AF interactions ($z=6,8$) that compares well with the FM case as $z$ increases. The key features of the specific heat curves, such as the shape of the peak and the high-temperature tail, are primarily determined by the effective number of neighbors and the single-ion anisotropy. These features are largely insensitive to the sign of the exchange interaction in the absence of strong frustration. For this reason we can also compare the experimental Néel temperature ($T_{N}$) with the theoretical Curie temperature ($T_{C}$).
@@
  • $|J_{3}/J_{1}|=3.7$ ($z=12$) which we can disregard has it seems unlikely to have a larger strength associated to a more distant bond.
  • $|J_{3}/J_{1}|=0.27$ ($z=12$) appears to be the only reasonable solution.
Without the actual simulation for this particular combination of couplings, we can only estimate the coupling strengths To do so, we use the mean-field result for the critical temperature (see section V.1).
@@
from which we obtain $|J_{1}|=|J_{2}|\sim 0.20$ K and $|J_{3}|\sim 0.05$ K. From these results we can not determine the signs of the coupling, but from the antiferromagnetic nature of the compound, one infers that at least one of these couplings should be antiferromagnetic.
GdPdBi adopts a half-Heusler structure with the Gd sublattice forming a face-centered cubic (FCC) lattice. Fitting with the layered honeycomb lattice (LHc(10), $z=10$) requires an unexpectedly large $J_{2}/J_{1}$ ratio of approximately $6.1$. To validate our approach, we compared our results with inelastic neutron scattering (INS) measurements on the related compound GdPtBi ($T_{N}=8.5$ K)[37, 38]. The LSWT analysis of GdPtBi[38] reported $J_{1}=0.14$ K, $J_{2}=0.27$ K, $J_{3}=0.049$ K, and $K=0.0098$ K, with $|J_{2}|/|J_{1}|>1$ (specifically $|J_{2}|/|J_{1}|\approx 1.93$), demonstrating that second-neighbor interactions dominate in this compound and that the anisotropy is small ($|K|/|J_{1}|\lesssim 0.1$). In GdPtBi, the magnetic structure consists of ferromagnetic (111) planes antiferromagnetically stacked along [111][37, 39], exhibiting low frustration. This structural and magnetic context validated the plausibility of the $|J_{2}|/|J_{1}|\gtrsim 1$ ratio for GdPdBi. But as shown in section VI the specific heat for $|J_{1}|<|J_{2}|$ is almost independent of $J_{2}$ and does not allow us to determine a ratio of interactions.
The crystal structure of GdNiSi3 is an orthorhombic system. Gd atoms forms a bilayered arrangement. Only interaction at the 7th nearest neighbor connects bilayers to provide a tridimensional connectivity ($z_{0}=17$). A good description of the experimental data is obtain considering up to 10th nearest-neighbor ($z=25$) while the best fit is seen considering up to 13th NN ($z=33$). This suggests that long range interactions, well beyond first neighbors, are relevant in this compound. Given the complexity of these interactions we only consider the uniform $J_{\delta}$ case. The GdNiSi3(33) lattice has $T_{C}/{|J|}\sim 145$ while experimentally $T_{N}=22$K so we can deduce $|J|\sim 22$K/145=0.15/145=0.15K. The large amount of couplings needed (up to 13 nearest-neighbors) certainly come from a distribution of interactions. This distribution could be determined with first principle calculations that considers at least this coordination.
@@
Now turning to Eu compounds, the importance of $K$ for these compounds may be attributed to the $\sim 25\%$ larger ionic radius of Eu2+ compared to Gd3+ [40], which makes the former more susceptible to crystal-field effects.
@@
The Eu sublattice of Eu2Pd3Sn can be though as a stacking of corrugated hexagonal layers in the ab plane. The smaller coordination number to have a tridimensional connectivity is $z_{0}=4$. This involves up to the third coordination shell of nearest neighbors (first and second nearest neighbors lie in the same layer). To properly compare the specific heat of Eu2Pd3Sn [41] with a Hamiltonian model we have to consider up to 6th nearest-neighbor in this lattice leading to a coordination $z=14$. The point group at the Eu sites is $C_{1h}$, which allows for the inclusion of the anisotropy term $H_{A}$, as discussed in section II.2. So an anisotropy term with $K/|J_{1}|=-0.4$ was included. With this anisotropy the Eu2Pd3Sn(14) lattice has $T_{C}/{|J_{0}|}\sim 66$. Experimentally $T_{N}=47$K so we can deduce, assuming uniform couplings, $|J|\sim 47$K/66=0.71/66=0.71K.
@@
Finally, the Eu sublattice of EuPdSn2 can be though as a stacking of rectangular layers in the ac plane. The smaller coordination number to have a tridimensional connectivity is $z_{0}=8$. This involves up to the third coordination shell of nearest-neighbor interactions. To properly compare the specific heat of EuPdSn2 [29] with a Hamiltonian model we have to consider up to 7th nearest-neighbor in this lattice leading to a coordination $z=22$. The point group at the Eu sites is $C_{2v}$, which also allows for the inclusion of the term $H_{A}$. So an anisotropy term with $K/|J_{0}|=-0.9$ was included. With this anisotropy the EuPdSn2 lattice has $T_{C}/{|J_{0}|}\sim 112$. Experimentally $T_{N}=12$K from which we deduce $|J|\sim 12$K/112=0.11/112=0.11K. The discrepancies near $T_{N}$ likely originate from factors not explicitly included in our model, such as more complex magnetic ordering (e.g., a collinear amplitude-modulated structure) resulting from modulated interactions and crystal-field effects, as discussed in Ref. 14.
@@
Since all the mentioned compounds contain transition-metal ions, which may be magnetic, one might question their influence on the specific heat. However, our analysis shows that the entropy obtained by integrating $C/T$ is close to $\ln(8)$ per rare-earth ion (consistent with $S=7/2$), indicating that any effects from transition metals are negligible.
@@
The success of these fits suggests that the effective number of neighbors $z$ is the dominant parameter determining the specific heat of 4f7 systems, with axial anisotropy $K$ providing additional refinement for Eu compounds. This approach significantly outperforms mean-field theory, which would fail to reproduce the observed specific heat curves (see Fig. 2).
@@
Figure 8: Experimental specific heat $C/T$ versus $T/T_{N}$ for six 4f7 compounds (dots) and our theoretical fits (lines). $K=0$ for all Gd-based compounds (GdNi3Ga9, GdCu2Ge2, GdPbBi, GdNiSi3). Inset for GdCu2Ge2 show $z$ as a function of $J_{3}/J_{1}$ (black line) and the fitting $z=12$ value (horizontal red line). Experimental data sources: Refs. 34, 35, 42, 36, 27, 29, 41.
VIII Summary